# Development & Characterization of Electrodes for large-scale Xenon Time Projection Chambers

### v1

@@ 1 Introduction
Time projection chambers (TPCs) are commonly used detectors in particle and astroparticle physics. They range from extremely large to rather compact scales. Large-scales TPCs are exemplified by the DUNE [DUNE_TPC, instruments5040031] liquid argon TPC which hosts 70,000 tonnes of target mass for neutrino science and the ALICE [ALME2010316] TPC with almost 90 m3 of a neon-based gas mixture for pattern recognition, tracking, and identification of charged particles in heavy-ion physics. Other more compact realizations are high-pressure gaseous xenon TPCs such as NEXT [next_epshep] with 100 kg target mass and the AXEL [HIKIDA2025170706] 180 L-size prototype, both used for searches for neutrinoless double beta decay. In direct Dark Matter (DM) searches, TPCs are the central element of multiple experiments. Among those is the DarkSide experiment, which featured a 50 kg liquid argon target mass, and will be expanding it to 20 tonnes [DS-50, DS-20k]. Furthermore, XENONnT [nt_instrument], LZ [lz_detector] and PandaX-4T [PandaX-4T-2019] are liquid-xenon (LXe) dual-phase TPCs with active masses of $\sim$ 6 tonnes (XENONnT, LZ) and $\sim$ 4 tonnes (PandaX-4T). Future TPCs in DM search are envisaged to grow even larger [Aalbers:2022dzr], with PandaX-20T reaching 20 tonnes with a TPC diameter of 2 m [PandaX-20T-2024], and XLZD with 60 to 80 tonnes of LXe target mass and a TPC measuring 3 m in diameter [XLZD:2024gxx].
@@
Electrodes are a central element of LXe dual-phase TPCs and their design and high-voltage performance becomes an increasing challenge with ever larger TPC dimensions. In this paper, after a general introduction into the working principle and design of electrodes (subsection 1.1 and subsection 1.2) we present a novel and comprehensive approach to design, build and test large-scale electrodes. We concentrate on two types of electrodes, a parallel-wire anode (section 2), and a hexagonal mesh cathode (section 3), both with a diameter of $\sim$$1.5\text{\,}\mathrm{m}$ typical for the XENONnT [nt_instrument] and LZ [lz_detector] experiments. In section 4, we report the high-voltage tests performed on the hexagonal mesh electrode as a way of quality assurance which can be generalized to other types of electrodes.
@@ 1.1 Signals in Dual-phase Xenon Time Projection Chambers
A typical xenon dual-phase TPC consists of a large cylindrical volume filled with liquid xenon (LXe) with a smaller gaseous layer (GXe) in the top part [nt_instrument, lz_detector, XLZD:2024gxx]. The top and bottom ends of the TPC are equipped with arrays of photomultiplier tubes (PMTs). Throughout the TPC, the electric fields are spanned by three electrodes: the negatively charged cathode at the bottom of the TPC and the gate electrode positioned right below the interface between liquid and gas span the so-called drift field $\vec{E}_{\text{drift}}$. The positively biased anode is situated in the gaseous phase and together with the gate defines the extraction and acceleration fields. In addition, in some configurations so-called screening electrodes are introduced in front of the PMT arrays to protect the sensors from electric fields [nt_instrument]. A diagram of a TPC can be seen in Figure 1.
Figure 1: Left: Diagram of a typical TPC [nt_instrument, lz_detector], showing the liquid and gaseous phases, the photo-sensors, the electrodes. In addition, the signal generation process is sketched out: An incident particle ($\chi$) scatters off Xe atoms, leading to an immediate scintillation signal (S1) and free electrons (e-) through ionization. The free electrons drift to the gaseous phase, where they are accelerated, generating a secondary scintillation pulse (S2). Right: Typical field strength defined by the electrodes, with the amplification, extraction, and drift field regions labeled.
@@
Correctly modeling the S2 signal requires a mostly homogeneous extraction field. Although small inhomogeneities can be corrected at the analysis level, abrupt changes caused, for example, by substructures on the electrodes are difficult to mitigate. Furthermore, a higher overall extraction field equals a higher chance of extracting the ionization electrons into the gas phase, increasing the strength of the S2 signal, and thus improving energy resolution and discrimination power between electronic and nuclear recoils. To ensure almost 100% extraction, a field in the gas near the surface of $E\sim$10\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$\$ is required [prop-scint]. Furthermore, the strength of the drift field must be optimized with respect to the recombination of electrons and Xe ions and the extent of drift within the TPC, seeking values of $E_{\text{drift}}\sim$290\text{\,}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$\$ [XLZD:2024gxx].
@@ 1.2 Design and Limitations of Electrodes for Time Projection Chambers
Commonly employed TPC designs use variations of parallel-wire electrodes [nt_instrument] or hexagonal mesh electrodes [next100_electrode, xenon100_instrument], or a combination of both [1t_instrument]. Additionally, there exist more novel designs such as woven-wire electrodes [lz_electrodes], or coated electrodes on a transparent substrate [darkside20k]. In this work, we focus on improving parallel-wire and hexagonal mesh electrodes. Their performance was studied in previous generations of experiments and showed significant limitations. Both designs have technical challenges to be solved for future detectors, in particular to guarantee minimal deformation, ease of installation and repair, and optimal manufacturing including quality control.
First, spanned parallel wires deform the ring preferentially in one direction, thus increasing the sagging and affecting the uniformity of the electric field. In contrast, a hexagonal mesh provides a more uniform tensioning and mechanical load in all directions on the plane, thus guaranteeing a more homogeneous bulk field between the electrodes. Additionally, regarding the ease of installing the electrode, the common strategy for a parallel-wire electrode with hundreds of wires is cumbersome and error-prone [lz_electrodes], unlike the case for an etched mesh, as demonstrated in [next100_electrode]. On the other hand, the etched mesh is monolithic and can hardly be repaired in case of imperfection after manufacturing or during handling.
@@
A further challenge for either type of electrode is the potential of spurious electron emission. Features, such as sharp edges, introduced during the production or handling of an electrode, can lead to localized extremely high fields. At such points, these field strengths can overcome the work function of the electrode material and release electrons into the detector [FN_field_emission, adam_bailey_thesis_em, PandaX:2014mem], in the worst case, causing electrical breakdown. As an example of structural deficiency, we report the case of the XENONnT detector. During its commissioning, a broken wire in one of its electrodes resulted in the inability to operate the TPC with a drift field greater than $\sim$$23\text{\,}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ [nt_instrument]. The detector remained functional and was successfully used for performing a variety of scientific searches [xenonnt_wimp_sr0, xenonnt_sr0_lower, xenonnt_b8]. However, without an upgrade of its electrodes to alleviate the underlying issue, the detector was not capable of reaching its design drift field and sensitivity [xenonnt_sensi].
@@ 2 Parallel-Wire Electrodes
Here we consider a scenario where a parallel-wire electrode is acting as an anode in a LXe dual-phase TPC. The developed electrode follows an initial design of the XENONnT experiment [nt_instrument, francesco_toschi_thesis] and consists of an array of stainless steel (SS) parallel wires. Each end of a wire was bent 90 individually and fixed to the SS electrode frame by a pin made from oxygen-free high conductivity (OFHC) copper ($99.99\%$ pure). The frame is a 24-gon with the inner diameter of $1.33\text{\,}\mathrm{m}$. The anode electrode is subject to several stringent requirements. To achieve field homogeneity, the separation between the wire centers is the same across the entire electrode, which is $5\text{\,}\mathrm{m}\mathrm{m}$, resulting in a total of 265 wires for the entire electrode. The gap between the anode and the liquid level being in the range from a few millimeters to a centimeter, significantly affects the amplification and width of the S2, leaving a small space for wire sagging. At the same time, the electrostatic force on the anode is about twice that on the gate due to the different dielectric constants of liquid and gaseous xenon. In addition, it is in the same direction as the gravitational force. Thus, the load on the anode wires is larger than that on the gate electrode.
To reduce sagging of the wires during detector operation, axial tension was applied to each wire during assembly. For long wires with small sagging, the load along the wire can be assumed to be uniform. Thus, we estimate the axial tensions $F_{ax}$ on the wires from the sagging model in [RoarkSS2022] :
@@
where $F_{ax}$ is the axial tension. $\omega_{g}$ and $\omega_{el}$ are the respective unit loads due to gravity and electrostatic force. $L_{0}$ is the undeformed length of the wire between supports. $h_{\text{max}}$ is the maximum sagging. The contribution from the elasticity of the wire was found to be less than $0.2\%$, and was therefore neglected [wire_with_elasticity, francesco_toschi_thesis]. We calculated $\omega_{el}$ using a 2D axially symmetric field simulation, considering a specific field configuration in the detector. Buoyancy is neglected in the load as it is four orders of magnitude smaller, given that the anode is in GXe. For a wire with a diameter of $0.216\text{\,}\mathrm{m}\mathrm{m}$, we estimated $(\omega_{g}+\omega_{el})$ to be $1.81\text{\times}{10}^{-5}\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}\mathrm{m}$ downwards. The undeformed length of the longest wire is $1340.4\text{\,}\mathrm{m}\mathrm{m}$, and the maximum sagging was targeted at $0.5\text{\,}\mathrm{m}\mathrm{m}$. Although the tension is proportional to $L^{2}$, we assort the wires into four groups of different tension values from 3.2 N per wire for the shortest ones up to 8.6 N per wire for the longest ones to ease the assembly process, as showcased in Figure 2. Due to the low tolerance on sagging and subsequently high axial tension required, we performed additional simulations and tests to investigate the structural integrity of the electrode.
@@ 2.1 Mechanical Simulations of the Electrode Frame
The frame is constructed from an annealed stainless steel SS316 slab with a yield tensile strength (YTS) of approximately $240\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$. Previous experience showed that the assembled electrode tended to relax after a cooling cycle, leading to increased sagging of the wires [francesco_toschi_thesis]. The annealed and mechanically relaxed frame was expected to reduce such effects.
@@
With iterations of the simulation, the geometry of the electrode frame was optimized to optimally withstand the stress of the wires. The height of the SS frame was increased to $24\text{\,}\mathrm{m}\mathrm{m}$ compared to the original XENONnT design of $18\text{\,}\mathrm{m}\mathrm{m}$, and the width varies along the circumference from 30.975 mm to 33.975 mm. The final maximum stress was 98.6 MPa, where the cross-section is the smallest. Compared to the YTS of the material, this design secured a safety factor of 2.4 against permanent plastic deformation. In addition, given the load of the wire tension, the ring should deform by a total of 4.41 mm outwards perpendicular to the wire direction. Parallel to the direction of the wires, the ring deformation is expected inwards by 4.44 mm in diameter. Based on this optimized design, the raw pre-manufactured ring was annealed, the final frame was then manufactured and electropolished by Mühlbauer Parts & System (MPS) [MPS].
@@ 2.2 Wire Tensile Tests at Different Temperatures
Table 1: Data sheet information for the candidates of the wire tests. YTS values refer to an offset of 0.2%. Measurements for UTS and YTS are presumably taken at room temperature.
@@
Vogelsang [vogelsang] 0.212 316 semi-hard 1000-1200 N.A.
Dahmen [Dahmen] 0.221 316L full-hard 1916.5 1528.5
@@
For each measurement, we first mounted a wire sample into the setup. The cold bath was prepared separately by adding dry ice to a bath of ethanol inside the partially thermally isolated bath. The liquid level was set to cover approximately half of the pulley. Two PT100 temperature sensors were used to monitor the temperature at different positions of the cold bath. As the temperature approached the desired value, we repositioned the entire L-shaped setup so that the horizontal part of the wire was completely submerged in the cold bath. A data acquisition module from National Instruments [national_instruments] continuously read out the data from the load cell and the temperature sensors during a measurement at the rate of $50\text{\,}\mathrm{H}\mathrm{z}$. The stretching screw was turned in steps of one-fourth of a turn while the stress was recorded by the load cell.
@@
To test the reliability of the pin fixation after a thermal cycle, we performed an additional tensile test after the system underwent a thermal cycle in a $200\text{\,}\mathrm{K}$ cold bath. We cooled down the setup for 1 minute until the measured temperature around the wire, monitored by a PT100 sensor, stabilized at $200\text{\,}\mathrm{K}$. Then, we took out the setup to warm it up until the temperature around the wire went back to room temperature before the tensile test. In case of a significant thermal relaxation between the copper pin fixation, the wire could slip, resulting in a reduction in the measured YTS and UTS values, which was not observed.
The resulting values of YTS for an offset of 0.01% at different temperatures are shown in Figure 4.
@@
The temperature was the average value recorded by the two PT100 sensors inside the bath. The error bound represents the maximum and minimum temperatures recorded by the sensors throughout each measurement. For each measurement, different fitting ranges that resulted in an $R^{2}$ value greater than 0.99 were used to obtain the YTS and the error. The length of the vertical part of the wire, which was not submerged in the cold bath, was checked to be an insignificant error compared to the fitting error. The error on the change in length and rolling friction was also estimated to be negligible. We validated the measurement by comparing the YTS at 0.2% offset to the data sheet values. For CFW, the mean value we measured at room temperature is consistent with the datasheet value, being $744\pm 12$ $\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ and $725\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ respectively. The error bar is the standard deviation of the measurements to account for systematics between configurations. For the Dahmen wire, the 0.2% offset at room temperature was out of the operating range of our setup. A datasheet value was not available for the Vogelsang wire sample.
@@
From our results, all wire samples tested can withstand the maximum stress, caused by the axial tension and the thermal stress, with less than 0.01% plastic deformation. Specifically, the maximum stress for the longest CFW wire is $11.35\text{\,}\mathrm{N}$, indicated by the dash-dotted line in Figure 3. Mechanical simulations indicate that under the gravitational and electrostatic forces, its sagging would subsequently be less than $0.5\text{\,}\mathrm{m}\mathrm{m}$. Compared to the estimated maximum stress, the average safety factor for CFW until deformation is $1.6\pm 0.2$. Additionally, as the temperature decreased to $200\text{\,}\mathrm{K}$, we observed a slight increasing trend in the YTS at a 0.01% offset, as suggested in [lz_electrodes], with correlation coefficients ranging from 0.62-0.62 to 0.68-0.68 for all wire candidates. The YTS at a 0.01% offset with copper pin fixation differs by less than 5% compared to the copper gasket fixation, regardless of whether the system underwent a thermal cycle.
@@
Apart from YTS and UTS, there are additional metrics to consider when selecting suitable candidates, namely the ductility and the surface treatment. Specifically, the Dahmen wire is prone to breaking during installation due to the stress from the copper pin fixation. Although the YTS is higher, hard wires are more brittle than other wires. Annealed SS, on the other hand, exhibits enhanced ductility among the three candidates, providing a longer range of plastic deformation before fracture. This property is ideal to avoid wire breakage, which could result in electrode short-circuiting or sparking at high voltages. Since CFW fulfills the mentioned requirements and has been tested in the XENONnT experiment, we subsequently used the CFW sample for the remaining tests.

### v2

@@ 1 Introduction
Time projection chambers (TPCs) are commonly used detectors in particle and astroparticle physics. They range from extremely large to rather compact scales. Large-scale TPCs are exemplified by the DUNE experiment [8, 7] which plans to operate a modular liquid argon TPC of 70,000 tonnes of total target mass for neutrino science and the ALICE [13] TPC with almost 90 m3 of a neon-based gas mixture for pattern recognition, tracking, and identification of charged particles in heavy-ion physics. Other more compact realizations are high-pressure gaseous xenon TPCs such as NEXT [58] with 100 kg target mass and the AXEL [47] 180 L-size prototype, both used for searches for neutrinoless double beta decay. In direct Dark Matter (DM) searches, TPCs are the central element of multiple experiments. Among those is the DarkSide experiment, which featured a 50 kg liquid argon target mass, and will be expanding it to 50 tonnes [9, 5]. Furthermore, XENONnT [25], LZ [12] and PandaX-4T [88] are liquid-xenon (LXe) dual-phase TPCs with active masses of $\sim$ 6 and $\sim$ 7 tonnes (XENONnT, LZ) and $\sim$ 4 tonnes (PandaX-4T). Future TPCs in DM searches are envisaged to grow even larger [1], with PandaX-xT reaching 43 tonnes with a TPC diameter of $\sim$ 2.6 m [6], and XLZD with 60 to 80 tonnes of LXe target mass and a TPC measuring 3 m in diameter [3].
@@
Electrodes are a central element of LXe dual-phase TPCs and their design and high-voltage performance becomes an increasing challenge with ever larger TPC dimensions. In this paper, after a general introduction into the working principle and design of electrodes (subsection 1.1 and subsection 1.2) we present a novel and comprehensive approach to design, build and test large-scale electrodes. We concentrate on two types of electrodes, a parallel-wire anode (section 2), and a hexagonal mesh cathode (section 3), both with a diameter of $\sim$$1.5\text{\,}\mathrm{m}$ typical for the XENONnT [25] and LZ [12] experiments. In section 4, we report the high-voltage tests performed on the hexagonal mesh electrode as a method of quality assurance which can be generalized to other types of electrodes.
@@ 1.1 Signals in Dual-phase Xenon Time Projection Chambers
A typical xenon dual-phase TPC consists of a large cylindrical volume filled with liquid xenon (LXe) with a smaller gaseous layer (GXe) in the top part [25, 12, 3]. The top and bottom ends of the TPC are equipped with arrays of photomultiplier tubes (PMTs). Throughout the TPC, the electric fields are spanned by three electrodes: the negatively charged cathode at the bottom of the TPC and the gate electrode positioned right below the interface between liquid and gas span the so-called drift field $\vec{E}_{\text{drift}}$. The positively biased anode is situated in the gaseous phase and together with the gate defines the extraction and acceleration fields. In addition, in some configurations so-called screening electrodes are introduced in front of the PMT arrays to protect the sensors from electric fields [25]. A diagram of a TPC can be seen in Figure 1.
Figure 1: Left: Diagram of a typical TPC [25, 12], showing the liquid and gaseous phases, the photo-sensors, the electrodes. In addition, the signal generation process is sketched out: An incident particle ($\chi$) scatters off Xe atoms, leading to an immediate scintillation signal (S1) and free electrons (e-) through ionization. The free electrons drift to the gaseous phase, where they are accelerated, generating a secondary scintillation pulse (S2). Right: Typical field strength defined by the electrodes, with the amplification, extraction, and drift field regions labeled.
@@
Correctly modeling the S2 signal requires a mostly homogeneous extraction field. Although small inhomogeneities can be corrected at the analysis level, abrupt changes caused, for example, by substructures on the electrodes are difficult to mitigate. Furthermore, a higher overall extraction field equals a higher chance of extracting the ionization electrons into the gas phase, increasing the strength of the S2 signal, and thus improving energy resolution and discrimination power between electronic and nuclear recoils. To ensure almost 100% extraction, a field in the gas near the surface of $E\sim$10\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$\$ is required [19]. Furthermore, the strength of the drift field must be optimized with respect to the recombination of electrons and Xe ions and the extent of drift within the TPC, seeking values of $E_{\text{drift}}\sim$290\text{\,}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$\$ [3].
@@ 1.2 Design and Limitations of Electrodes for Time Projection Chambers
Commonly employed TPC designs use variations of parallel-wire electrodes [25] or hexagonal mesh electrodes [60, 20], or a combination of both [21]. Additionally, there exist more novel designs such as woven-wire electrodes [56], or coated electrodes on a transparent substrate [4]. In this work, we focus on improving parallel-wire and hexagonal mesh electrodes. Their performance was studied in previous generations of experiments and showed significant limitations. Both designs have technical challenges to be solved for future detectors, in particular to guarantee minimal deformation, ease of installation and repair, and optimal manufacturing including quality control.
First, spanned parallel wires deform the ring preferentially in one direction, thus increasing the sagging and affecting the uniformity of the electric field. In contrast, a hexagonal mesh provides a more uniform tensioning and mechanical load in all directions on the plane, thus guaranteeing a more homogeneous bulk field between the electrodes. Additionally, regarding the ease of installing the electrode, the common strategy for a parallel-wire electrode with hundreds of wires is cumbersome and error-prone [56], unlike the case for an etched mesh, as demonstrated in [60]. On the other hand, the etched mesh is monolithic and can hardly be repaired in case of imperfection after manufacturing or during handling.
@@
A further challenge for either type of electrode is the potential of spurious electron emission. Features, such as sharp edges, introduced during the production or handling of an electrode, can lead to localized extremely high fields. At such points, these field strengths can overcome the work function of the electrode material and release electrons into the detector [42, 28, 34], in the worst case, causing electrical breakdown. As an example of structural deficiency, we report the case of the XENONnT detector. During its commissioning, a broken wire in one of its electrodes resulted in the inability to operate the TPC with a drift field greater than $\sim$$23\text{\,}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ [25]. The detector remained functional and was successfully used for performing a variety of scientific searches [24, 23, 16]. However, without an upgrade of its electrodes to alleviate the underlying issue, the detector was not capable of reaching its design drift field and sensitivity [17].
@@ 2 Parallel-Wire Electrodes
Here we consider a scenario where a parallel-wire electrode is acting as an anode in a LXe dual-phase TPC. The developed electrode follows an initial design of the XENONnT experiment [25, 81] and consists of an array of stainless steel (SS) parallel wires. Each end of a wire was bent 90 individually and fixed to the SS electrode frame by a pin made from oxygen-free high conductivity (OFHC) copper ($99.99\%$ pure). The frame is a 24-gon with the inner diameter of $1.33\text{\,}\mathrm{m}$. The anode electrode is subject to several stringent requirements. To achieve field homogeneity, the separation between the wire centers is the same across the entire electrode, which is $5\text{\,}\mathrm{m}\mathrm{m}$, resulting in a total of 265 wires for the entire electrode. The gap between the anode and the liquid level being in the range from a few millimeters to a centimeter, significantly affects the amplification and width of the S2, leaving a small space for wire sagging. At the same time, the electrostatic force on the anode is about twice that on the gate due to the different dielectric constants of liquid and gaseous xenon. In addition, it is in the same direction as the gravitational force. Thus, the load on the anode wires is larger than that on the gate electrode. The requirements for a parallel-wire electrode as an anode are listed in Table 1.
Table 1: Requirements for a parallel-wire electrode as an anode. The parameter AA is the cross-sectional area of the wire. The thermal stress ($\sigma_{s}$) was calculated at $75\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ assuming the Young’s modules of the wire to be $200\text{\,}\mathrm{G}\mathrm{P}\mathrm{a}$ at $175\text{\,}\mathrm{K}$ [64], the coefficient of thermal expansion at $1.5\text{\times}{10}^{-5}\text{\,}$ [64], and the temperature difference across the electrode at $25\text{\,}\mathrm{K}$. The yield tensile strength ($\text{YTS}_{0.01\%,\,\text{meas.}}$) was measured as described in subsection $2.$ The definition of the showcased parameters $h_{\text{max}}$ and $F_{ax}$ is provided in the text below.
Quantity Symbol Requirement
Frame deformation $5\text{\,}\mathrm{m}\mathrm{m}$
Sagging limit $h_{\text{max}}$ $0.5\text{\,}\mathrm{m}\mathrm{m}$
Safety factor $\text{YTS}_{0.01\%,\,\text{meas.}}\,/\,(\sigma_{s}+F_{ax}/A)$ $\gtrsim 1.5$
Transparency $\gtrsim 90\%$
To reduce sagging of the wires during detector operation, axial tension was applied to each wire during assembly. For long wires with small sagging, the load along the wire can be assumed to be uniform. Thus, we estimate the axial tensions $F_{ax}$ on the wires from the sagging model in [87] :
@@
where $F_{ax}$ is the axial tension. $\omega_{g}$ and $\omega_{el}$ are the respective unit loads due to gravity and electrostatic force. $L_{0}$ is the undeformed length of the wire between supports. $h_{\text{max}}$ is the maximum sagging. The contribution from the elasticity of the wire was found to be less than $0.2\%$, and was therefore neglected [49, 81]. We calculated $\omega_{el}$ using a 2D axially symmetric field simulation, considering a specific field configuration in the detector. Buoyancy is neglected in the load as it is four orders of magnitude smaller, given that the anode is in GXe. For a wire with a diameter of $0.216\text{\,}\mathrm{m}\mathrm{m}$, we estimated $(\omega_{g}+\omega_{el})$ to be $1.81\text{\times}{10}^{-5}\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}\mathrm{m}$ downwards. The undeformed length of the longest wire is $1340.4\text{\,}\mathrm{m}\mathrm{m}$, and the maximum sagging was targeted at $0.5\text{\,}\mathrm{m}\mathrm{m}$. Although the tension is proportional to $L_{0}^{2}$, we assort the wires into four groups of different tension values from 3.2 N per wire for the shortest ones up to 8.6 N per wire for the longest ones to ease the assembly process, as showcased in Figure 2. Due to the low tolerance on sagging and subsequently high axial tension required, we performed additional simulations and tests to investigate the structural integrity of the electrode.
@@ 2.1 Mechanical Simulations of the Electrode Frame
The frame is constructed from a stainless steel SS316 slab with a yield tensile strength (YTS) of approximately $240\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ [38]. Previous experience showed that the assembled electrode tended to release internal stress after a cooling cycle, leading to increased sagging of the wires [81]. To avoid such relaxation of the frame, it was annealed before the final machining to relax its internal mechanical stress.
@@
With iterations of the simulation, the geometry of the electrode frame was optimized to optimally withstand the stress of the wires. The height of the SS frame was increased to $24\text{\,}\mathrm{m}\mathrm{m}$ compared to the original XENONnT design of $18\text{\,}\mathrm{m}\mathrm{m}$, and the width varies along the circumference from 30.975 mm to 33.975 mm. The final maximum stress was 98.6 MPa, where the cross-section is the smallest. Compared to the YTS of the material, this design secured a safety factor of 2.4 against permanent plastic deformation. In addition, given the load of the wire tension, the ring should deform by a total of 4.41 mm outwards perpendicular to the wire direction. Parallel to the direction of the wires, the ring deformation is expected inwards by 4.44 mm in diameter. Based on this optimized design, the raw pre-manufactured ring was annealed, the final frame was then manufactured and electropolished by Mühlbauer Parts & System (MPS) [63].
@@ 2.2 Wire Tensile Tests at Different Temperatures
Table 2: Data sheet information for the candidates of the wire tests. YTS values refer to an offset of 0.2%. Measurements for ultimate tensile strength (UTS) and YTS are presumably taken at room temperature.
@@
Vogelsang [84] 0.212 316 semi-hard 1000-1200 N.A.
Dahmen [50] 0.221 316L full-hard 1916.5 1528.5
@@
For each measurement, we first mounted a wire sample into the setup. The cold bath was prepared separately by adding dry ice to a bath of ethanol inside the partially thermally isolated bath. The liquid level was set to cover approximately half of the pulley. Two PT100 temperature sensors were used to monitor the temperature at different positions of the cold bath. As the temperature approached the desired value, we repositioned the entire L-shaped setup so that the horizontal part of the wire was completely submerged in the cold bath. A data acquisition module from National Instruments [65] continuously read out the data from the load cell and the temperature sensors during a measurement at the rate of $50\text{\,}\mathrm{H}\mathrm{z}$. The stretching screw was turned in steps of one-fourth of a turn while the stress was recorded by the load cell.
@@
To test the reliability of the pin fixation after a thermal cycle, we performed an additional tensile test after the system underwent a thermal cycle in a $200\text{\,}\mathrm{K}$ cold bath, which is close to the detector operating temperature of $175\text{\,}\mathrm{K}$ [25]. We cooled down the setup for 1 minute until the measured temperature around the wire, monitored by a PT100 sensor, stabilized at $200\text{\,}\mathrm{K}$. Then, we took out the setup to warm it up until the temperature around the wire went back to room temperature before the tensile test. In case of a significant thermal relaxation between the copper pin fixation, the wire could slip, resulting in a reduction in the measured YTS and ultimate tensile strength (UTS) values, which was not observed.
The resulting values of YTS for an offset of 0.01% as well as the UTS at different temperatures are shown in Figure 4.
@@
The temperature was the average value recorded by the two PT100 sensors inside the bath. The error bound represents the maximum and minimum temperatures recorded by the sensors throughout each measurement. For each measurement, the YTS and its error were obtained from linear regressions performed over different fitting ranges, with the ranges chosen to ensure $R^{2}>0.99$. The length of the vertical part of the wire, which was not submerged in the cold bath, was checked to be an insignificant error compared to the fitting error. The error on the change in length and rolling friction was also estimated to be negligible. We validated the measurement by comparing the YTS at 0.2% offset to the data sheet values. For CFW, the mean value we measured at room temperature is consistent with the datasheet value, being $744\pm 12$ $\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ and $725\text{\,}\mathrm{M}\mathrm{P}\mathrm{a}$ respectively. The error bar is the standard deviation of the measurements to account for systematics between configurations. For the Dahmen wire, the 0.2% offset at room temperature was out of the operating range of our setup. A datasheet value was not available for the Vogelsang wire sample. The UTS was determined as the maximum stress in the stress–strain curve for each measurement, provided that the wire sample broke. Similarly to the YTS measurement, for some wire samples from Dahmen and Volgesang, breakage did not occur within the maximum strain achievable with the experimental setup.
@@
From our results, all wire samples tested can withstand the maximum stress, caused by the axial tension and the thermal stress, with less than 0.01% plastic deformation. Specifically, the maximum stress for the longest CFW wire is $11.35\text{\,}\mathrm{N}$, indicated by the dash-dotted line in Figure 3. Mechanical simulations indicate that under the gravitational and electrostatic forces, its sagging would subsequently be less than $0.5\text{\,}\mathrm{m}\mathrm{m}$. Compared to the estimated maximum stress, the average safety factor for CFW until deformation is $1.6\pm 0.2$. Additionally, as the temperature decreased to $200\text{\,}\mathrm{K}$, we observed a slight increasing trend in the YTS at a 0.01% offset, as suggested in [56], with correlation coefficients ranging from 0.62-0.62 to 0.68-0.68 for all wire candidates. The YTS at a 0.01% offset with copper pin fixation differs by less than 5% compared to the copper gasket fixation, regardless of whether the system underwent a thermal cycle.
@@
Apart from YTS and UTS, there are additional metrics to consider when selecting suitable candidates, namely the ductility and the surface treatment. Specifically, the Dahmen wire is prone to breaking during installation due to the stress from the copper pin fixation. Although the YTS is higher, hard wires are more brittle than other wires. Annealed SS, on the other hand, exhibits enhanced ductility among the three candidates, providing a longer range of plastic deformation before fracture. This property is ideal to avoid wire breakage, which could result in electrode short-circuiting or sparking at high voltages. Since CFW fulfills the mentioned requirements and has been tested in the XENONnT experiment, we subsequently used the CFW sample for the remaining tests. With a pitch of 5 mm and the wire cross section of 0.216 mm, an optical transparency of 95.7% is obtained, fulfilling the requirement stated in Table 1.